# End-to-End Fine-Tuning of 3D Texture Generation using Differentiable Rewards

### v2

@@ 3.2 Texture Rewards Learning
where $r$ is a differentiable reward function, $\operatorname{TexGen}(\boldsymbol{\theta},\mathbf{c},\mathbf{v_{gen}})$ represents the iterative texturing process from different viewpoints $v_{gen}$ (explained in Sec. 3.1 and Appendix A), and $c$ is the text prompt randomly sampled from the text dataset $p_{c}$. To solve Eq. 1, instead of maximizing the reward function, we consider the negative of it as a loss function $L(\theta)=-J(\theta)$. Then, we compute $-\nabla_{\boldsymbol{\theta}}r\left(\operatorname{TexGen}\left(\boldsymbol{\theta},\mathbf{c},\mathbf{v_{gen}}\right),\mathbf{c}\right)$ and update $\theta$ using gradient-based optimization algorithms and the following update rule: $\theta\leftarrow\theta+\eta\nabla_{\boldsymbol{\theta}}r\left(\operatorname{TexGen}\left(\boldsymbol{\theta},\mathbf{c},\mathbf{v_{gen}}\right),\mathbf{c}\right)$. However, naively optimizing the $\theta$ would results in requiring a massive amount of memory which makes the fine-tuning process impractical. This is because of the fact that computing the gradients $\nabla_{\boldsymbol{\theta}}r$ requires backpropagation through multiple diffusion steps and multiple camera viewpoints which stems from the texturing process $\operatorname{TexGen}(\boldsymbol{\theta},\mathbf{c},\mathbf{v_{gen}})$. Consequently, this necessitates storing all the intermediate variables during the multi-view texturing process. More specifically, each step in the texturing process involves at least 10 diffusion steps, and with 8–10 total texturing steps (corresponding to the number of camera viewpoints), this results in approximately 100 diffusion steps overall. Each step requires around 3 GB of GPU memory, leading to a total memory demand of roughly 300 GB to train the texture-generation pipeline. Such hardware requirements are nearly infeasible, especially in typical academic environments. To alleviate this issue, we leverage two main approaches to reduce the memory used by our method during the training stage: 1) low-rank adaptation (LoRA) [13], and 2) gradient checkpointing [5]. Applying LoRA to the U-Net in the latent diffusion model reduces trainable parameters by roughly a factor of 1000 compared to full fine-tuning. Additionally, two levels of activation checkpointing keep memory usage constant across views and diffusion steps, enabling higher-resolution renders and more diffusion steps. Further memory-saving details are provided in Appendix B.
@@ 3.3 Geometry-Aware Reward Design
4:  Initialize 3D asset parameters $a$: texture $albedo$, weight $cnt$, view cache $viewcos$
@@
[…]
@@
$\mathbf{d}_{i}^{\mathbf{uv}}=\frac{\mathbf{u}_{\text{end }}-\mathbf{u}_{\text{orig }}}{\left\|\mathbf{u}_{\text{end }}-\mathbf{u}_{\text{orig }}\right\|}$ (4)
@@
(III) Projecting 3D mirror points into UV space. To obtain the UV coordinates corresponding to $\mathbf{v}_{i}^{\text{mirror }}$, we perform a closest-point query on the mesh surface to find the nearest triangle containing the surface projection of $\mathbf{v}_{i}^{\text{mirror }}$. Given the triangle’s 3D vertices $\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3}$, and corresponding UV coordinates $\mathbf{u}_{1},\mathbf{u}_{2},\mathbf{u}_{3}\in[0,1]^{2}$, we compute barycentric coordinates $(\beta_{1},\beta_{2},\beta_{3})$ of the projected point and obtain the UV coordinates of the mirror as: $\mathbf{u}_{i}^{\text{mirror }}=\beta_{1}\mathbf{u}_{1}+\beta_{2}\mathbf{u}_{2}+\beta_{3}\mathbf{u}_{3}$. We store the UV coordinate $\mathbf{u}_{i}$ of the original vertex along with $\mathbf{u}_{i}^{\text{mirror}}$ and form a pair $(\mathbf{u}_{i},\mathbf{u}_{i}^{\text{mirror}})$ used in our symmetry-awared reward.
@@
where $r$ denotes the LAION aesthetic reward [29], and $\operatorname{TexGen}(\boldsymbol{\theta},\mathbf{c},\mathbf{v}{\mathrm{gen}}(\boldsymbol{\phi}))$ corresponds to the texture‐generation step of Sec. 3.1, with the diffusion model parameters $\boldsymbol{\theta}$ held fixed and the camera viewpoints $\mathbf{v}_{\mathrm{gen}}(\boldsymbol{\phi})$ learned instead of predefined. To solve Eq. 10, instead of maximizing the aesthetic reward, we consider the negative of it as a loss function $L(\phi)=-J(\phi)$. Then, we compute $-\nabla_{\boldsymbol{\phi}}r\left(\operatorname{TexGen}\left(\boldsymbol{\theta},\mathbf{c},\mathbf{v_{gen}(\phi)}\right),\mathbf{c}\right)$ and update $\phi$ using gradient-based optimization algorithms and the following update rule: $\phi\leftarrow\phi+\eta\nabla_{\boldsymbol{\phi}}r\left(\operatorname{TexGen}\left(\boldsymbol{\theta},\mathbf{c},\mathbf{v_{gen}(\phi)}\right),\mathbf{c}\right)$.
@@ A.2 Modifications To Make Texture Generation Step Differentiable
$\begin{array}[]{lr}\mathbf{f}=\text{ normalize }(\mathbf{c}-\mathbf{t})&\text{ (forward vector) },\\ \mathbf{r}=\text{ normalize }(\mathbf{up}\times\mathbf{f})&\text{ (right vector) },\\ \mathbf{u}=\text{ normalize }(\mathbf{f}\times\mathbf{r})&\text{ (up vector) },\end{array}$ (17)
@@
$M_{\text{keep }}$ =$=\neg M_{\text{generate }}\cdot\neg M_{\text{refine }}$
where 1[]1[$\cdot$]1 [ ⋅ ] denotes the indicator function and logical operators like, $\neg$ produce discrete masks, making them non-differentiable and unsuitable for gradient-based optimization. To enable end-to-end differentiability, we replace these hard decisions with soft, continuous approximations using the sigmoid function. Let $\sigma_{k}(x)=\frac{1}{1+\exp(-kx)}$ denote the sigmoid function with steepness $k$. We define soft masks as follows: (i) soft generation mask -
@@ Appendix C Remarks on Principal Curvature Directions
An interesting observation in our results (see Fig. 3 and Fig. 10) is the emergence of repetitive texture patterns after fine-tuning with the geometry-texture alignment reward. This behavior stems from the mismatch between the number of curvature vectors and the number of texture gradients. Specifically, curvature is a vertex-level attribute defined per mesh vertex, whereas texture gradients are defined per pixel in the 2D UV space. For example, in the bunny mesh example, we have 4487 curvature vectors (one per vertex), but over one million texture gradients from a $\times 1024$ 1024 texture image. To compute cosine similarity between curvature vectors and texture gradients, both must be in the same space and of the same dimension. We address this by sampling the texture gradients at the UV coordinates corresponding to mesh vertices, the same positions where curvature vectors are defined. This results in two aligned tensors (both of size 4487 in the bunny case), enabling the reward function to compare directional alignment directly. Over time, this differentiable sampling mechanism guides the model to place strong texture features (e.g., edges) at those specific UV coordinates. As these positions remain fixed (being tied to the mesh vertices), the model learns to reinforce edge features at the same UV locations across views and iterations. This alignment results in visually repetitive patterns in the generated textures, especially in regions of high curvature.
@@ Appendix E Remarks on Texture Features Emphasis Reward
[…]
@@
where $\sigma_{rg}$, $\sigma_{yb}$, $\mu_{rg}$, and $\mu_{yb}$ are the standard deviations and mean colors offsets corresponding to color opponent components $rg(x,y)=R(x,y)-G(x,y)$ and $yb(x,y)=\frac{R(x,y)+G(x,y)}{2}-B(x,y)$, respectively. The full texture features emphasis reward combines both components:
@@ F.1 Geometry-Aware Texture Colorization Reward Design
[…]
@@
$\operatorname{val}=$ $\operatorname{bary}[0]\cdot\operatorname{curv}_{v}[\operatorname{tri\_v}[0]]+$
@@ Appendix G Attempts to Overcome Reward Hacking
LoRA scaling. Inspired by [7] and similar to [31], we scale the LoRA parameters by a factor in the range (0,1](0,1]( 0 , 1 ] to modulate their effect during fine-tuning. Empirically, we observe that factors close to 1, specifically within [0.9,1][0.9,1][ 0.9 , 1 ], consistently yield the best performance in our experiments.
@@
where $\sigma_{rg}$, $\sigma_{yb}$, $\mu_{rg}$, and $\mu_{yb}$ are the standard deviations and mean colors offsets corresponding to color opponent components $rg(x,y)=R(x,y)-G(x,y)$ and $yb(x,y)=\frac{R(x,y)+G(x,y)}{2}-B(x,y)$, respectively. This term is incorporated into both our texture feature emphasis and symmetry-aware texture generation reward functions to promote visually rich and saturated outputs.

### v3

@@ 3.2 Texture Rewards Learning
where $r$ is a differentiable reward function, $\operatorname{TexGen}(\boldsymbol{\theta},\mathbf{c},\mathbf{v_{gen}})$ represents the iterative texturing process from different viewpoints $v_{gen}$ (explained in Sec. 3.1 and Appendix A), and $c$ is the text prompt randomly sampled from the text dataset $p_{c}$. To solve Eq. 1, instead of maximizing the reward function, we consider the negative of it as a loss function $L(\theta)=-J(\theta)$. Then, we compute $\nabla_{\boldsymbol{\theta}}r\left(\operatorname{TexGen}\left(\boldsymbol{\theta},\mathbf{c},\mathbf{v_{gen}}\right),\mathbf{c}\right)$ and update $\theta$ using gradient-based optimization algorithms and the following update rule: $\theta\leftarrow\theta+\eta\nabla_{\boldsymbol{\theta}}r\left(\operatorname{TexGen}\left(\boldsymbol{\theta},\mathbf{c},\mathbf{v_{gen}}\right),\mathbf{c}\right)$. However, naively optimizing the $\theta$ would results in requiring a massive amount of memory which makes the fine-tuning process impractical. This is because of the fact that computing the gradients $\nabla_{\boldsymbol{\theta}}r$ requires backpropagation through multiple diffusion steps and multiple camera viewpoints which stems from the texturing process $\operatorname{TexGen}(\boldsymbol{\theta},\mathbf{c},\mathbf{v_{gen}})$. Consequently, this necessitates storing all the intermediate variables during the multi-view texturing process. More specifically, each step in the texturing process involves at least 10 diffusion steps, and with 8–10 total texturing steps (corresponding to the number of camera viewpoints), this results in approximately 100 diffusion steps overall. Each step requires around 3 GB of GPU memory, leading to a total memory demand of roughly 300 GB to train the texture-generation pipeline. Such hardware requirements are nearly infeasible, especially in typical academic environments. To alleviate this issue, we leverage two main approaches to reduce the memory used by our method during the training stage: 1) low-rank adaptation (LoRA) [13], and 2) gradient checkpointing [5]. Applying LoRA to the U-Net in the latent diffusion model reduces trainable parameters by roughly a factor of 1000 compared to full fine-tuning. Additionally, two levels of activation checkpointing keep memory usage constant across views and diffusion steps, enabling higher-resolution renders and more diffusion steps. Further memory-saving details are provided in Appendix B.
@@ 3.3 Geometry-Aware Reward Design
4:  Initialize 3D asset parameters $a$: texture albedoalbedo, weight cntcnt, view cache viewcosviewcos
@@
Texture Features Emphasis Reward. This reward encourages texture patterns that emphasize 3D surface structure while maintaining perceptual richness through color variation. Specifically, it strengthens texture variations in high-curvature regions to enhance capturing geometric structures. Let $T(x,y)\in[0,1]^{3}$ denote the RGB texture at pixel $(x,y)$ and $C(x,y)\in[0,1]$ be the normalized curvature map in UV space. The texture-curvature alignment term is mathematically defined as the negative mean squared error (MSE):
@@
[…]
@@
[…]
@@
where $r$ denotes the LAION aesthetic reward [29], and $\operatorname{TexGen}(\boldsymbol{\theta},\mathbf{c},\mathbf{v}{\mathrm{gen}}(\boldsymbol{\phi}))$ corresponds to the texture‐generation step of Sec. 3.1, with the diffusion model parameters $\boldsymbol{\theta}$ held fixed and the camera viewpoints $\mathbf{v}_{\mathrm{gen}}(\boldsymbol{\phi})$ learned instead of predefined. To solve Eq. 10, instead of maximizing the aesthetic reward, we consider the negative of it as a loss function $L(\phi)=-J(\phi)$. Then, we compute $\nabla_{\boldsymbol{\phi}}r\left(\operatorname{TexGen}\left(\boldsymbol{\theta},\mathbf{c},\mathbf{v_{gen}(\phi)}\right),\mathbf{c}\right)$ and update $\phi$ using gradient-based optimization algorithms and the following update rule: $\phi\leftarrow\phi+\eta\nabla_{\boldsymbol{\phi}}r\left(\operatorname{TexGen}\left(\boldsymbol{\theta},\mathbf{c},\mathbf{v_{gen}(\phi)}\right),\mathbf{c}\right)$.
@@ 5 Conclusions
Acknowledgements
We acknowledge funding from the Mila Tech Transfer Grant with Silicolabs and internal funds Concordia University. Computational resources were provided by Calcul Québec and the Digital Research Alliance of Canada.
@@ A.2 Modifications To Make Texture Generation Step Differentiable
[…]
@@
$M_{\text{keep }}$ $=\neg M_{\text{generate }}\cdot\neg M_{\text{refine }}$
where $1[\cdot]$ denotes the indicator function and logical operators like, $\neg$ produce discrete masks, making them non-differentiable and unsuitable for gradient-based optimization. To enable end-to-end differentiability, we replace these hard decisions with soft, continuous approximations using the sigmoid function. Let $\sigma_{k}(x)=\frac{1}{1+\exp(-kx)}$ denote the sigmoid function with steepness $k$. We define soft masks as follows: (i) soft generation mask -
@@ Appendix C Remarks on Principal Curvature Directions
An interesting observation in our results (see Fig. 3 and Fig. 10) is the emergence of repetitive texture patterns after fine-tuning with the geometry-texture alignment reward. This behavior stems from the mismatch between the number of curvature vectors and the number of texture gradients. Specifically, curvature is a vertex-level attribute defined per mesh vertex, whereas texture gradients are defined per pixel in the 2D UV space. For example, in the bunny mesh example, we have 4487 curvature vectors (one per vertex), but over one million texture gradients from a $1024\times 1024$ texture image. To compute cosine similarity between curvature vectors and texture gradients, both must be in the same space and of the same dimension. We address this by sampling the texture gradients at the UV coordinates corresponding to mesh vertices, the same positions where curvature vectors are defined. This results in two aligned tensors (both of size 4487 in the bunny case), enabling the reward function to compare directional alignment directly. Over time, this differentiable sampling mechanism guides the model to place strong texture features (e.g., edges) at those specific UV coordinates. As these positions remain fixed (being tied to the mesh vertices), the model learns to reinforce edge features at the same UV locations across views and iterations. This alignment results in visually repetitive patterns in the generated textures, especially in regions of high curvature.
@@ Appendix E Remarks on Texture Features Emphasis Reward
This reward encourages texture patterns that emphasize 3D surface structure while maintaining perceptual richness through color variation. It consists of two components: (i) texture-curvature magnitude alignment which strengthens texture variations in high-curvature regions to enhance capturing geometric structures, and (ii) colorfulness inspired by [10, 12], which discourages desaturated textures by encouraging vibrant color use, measured through the standard deviation and mean distance of opponent color channels from neutral gray. Let $T(x,y)\in[0,1]^{3}$ denote the RGB texture at pixel $(x,y)$ and $C(x,y)\in[0,1]$ be the normalized curvature map in UV space. The texture-curvature alignment term is mathematically defined as the negative mean squared error (MSE):
@@
where $\sigma_{rg}$, $\sigma_{yb}$, $\mu_{rg}$, and $\mu_{yb}$ are the standard deviations and mean colors offsets corresponding to color opponent components rg(x,y)=R(x,y)G(x,y)rg(x,y)=R(x,y)-G(x,y) and $yb(x,y)=\frac{R(x,y)+G(x,y)}{2}-B(x,y)$, respectively. The full texture features emphasis reward combines both components:
@@ F.1 Geometry-Aware Texture Colorization Reward Design
To encode surface curvature into texture color, we design a reward function that encourages warm colors (e.g., red, yellow) in regions of high curvature and cool colors (e.g., blue, green) in regions of low curvature. Given a per-pixel scalar curvature map $C(x,y)\in[-1,1]$, and a curvature threshold $T\in[-1,1]$, let $I_{r}(x,y)$ and $I_{b}(x,y)$ denote the red and blue channels of the predicted texture at pixel $(x,y)$, respectively. For pixels where $C(x,y)>T$, we encourage warm colors by rewarding a higher red-blue channel difference $\Delta_{rb}(x,y)=I_{r}(x,y)-I_{b}(x,y)$. Conversely, for $C(x,y)<T$, we encourage cool colors by penalizing this difference. The final reward is the average $\Delta_{rb}$ across all pixels:
@@
$\operatorname{val}=\$$\operatorname{bary}[0]\cdot\operatorname{curv}_{v}[\operatorname{tri\_v}[0]]+$
@@ Appendix G Attempts to Overcome Reward Hacking
LoRA scaling. Inspired by [7] and similar to [31], we scale the LoRA parameters by a factor in the range $(0,1]$ to modulate their effect during fine-tuning. Empirically, we observe that factors close to 1, specifically within $[0.9,1]$, consistently yield the best performance in our experiments.
@@
where $\sigma_{rg}$, $\sigma_{yb}$, $\mu_{rg}$, and $\mu_{yb}$ are the standard deviations and mean colors offsets corresponding to color opponent components rg(x,y)=R(x,y)G(x,y)rg(x,y)=R(x,y)-G(x,y) and $yb(x,y)=\frac{R(x,y)+G(x,y)}{2}-B(x,y)$, respectively. This term is incorporated into both our texture feature emphasis and symmetry-aware texture generation reward functions to promote visually rich and saturated outputs.